Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Orlati Andrea at Fri Feb 15 05:43:11 MET 2002) Success
Experiment successfully completed at 4.20 UT. Nothing to signal.

Torun (Kaz Borkowski at Fri Feb 15 09:09:39 MET 2002) Success
No known problems (remarks for the wavelength do not apply for now Tsys measurements are back and there were only 4 BBCs in use).

Onsala (John Conway at Fri Feb 15 09:27:58 MET 2002) Success
Occasional RFI for short periods otherwise no known problems.

Effelsberg (A. Kraus at Fri Feb 15 13:53:06 MET 2002) Success

Jodrell_Bank (Alastair Gunn at Mon Feb 18 11:05:31 MET 2002) Minor failures
Several minor drive motor dropouts noted in log. No other known problems.

Cambridge (Alastair Gunn at Mon Feb 18 11:06:41 MET 2002) Minor failures
Cambridge telescope back online so ran schedule. However, recorder has developed a serious fault and will be taken offline after this experiment. Impossible to indicate the quality of the data. No other problems noticed.

Onsala (John Conway at Tue Feb 19 08:35:36 MET 2002) Minor failures
No significant problems.

Westerbork (Tony Foley at Tue Feb 19 14:29:43 MET 2002) Minor failures
One scan at 23:08 out of HA range. Beware possible 180deg phase jumps about 17:22, 20:57, 22:28, 00:25, 00:32, 00:38, 02:37, 02:56, 03:38, 03:46 and 04:15UT

Urumqi (Liu Xiang at Fri Mar 22 14:02:49 CET 2002) Success
except many sources too low, noknown problems, but parity error, see log file in detail.